Kyky232:

In triangles GHI and RST, ∠G ≅ ∠R, ∠H ≅ ∠S, and segment GI ≅ segment RT. Is this information sufficient to prove triangles GHI and RST congruent through SSS? Explain your answer.

jhonyy9:

 so from these given what result - how you think ? look pls on my posted image

Kyky232:

Side Side Side postulate would mean that they have the same sides? Both do?

Kyky232:

I would say no, right?

jhonyy9:

ok pls step by step what is the sum of angles measure of a triangle ?

Kyky232:

180?

jhonyy9:

exactly so than we have that from 3 angles measure 2 are the same - so equal how will be the 3rd ?

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 exactly so than we have that from 3 angles measure 2 are the same - so equal how will be the 3rd ? \(\color{#0cbb34}{\text{End of Quote}}\) Now Im confused. Sorry Im slow...

xxDeppressionxx:

No one is slow they just don't comprehend what people say sometimes

jhonyy9:

ok there are angles A and B and C on the second triangle D and E and F so than given that A = D B = E how are C and F ?

jhonyy9:

so are equal too

jhonyy9:

bc. 180 -(A+B) = 180 - (D+E) yes ?

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 bc. 180 -(A+B) = 180 - (D+E) yes ? \(\color{#0cbb34}{\text{End of Quote}}\) yeah?

Kyky232:

I think?

braids36:

h2o

Timmyspu:

I can confirm that @jhonyy9 is right

gabrielchestnut:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Timmyspu I can confirm that @jhonyy9 is right \(\color{#0cbb34}{\text{End of Quote}}\) me too

jhonyy9:

ok so than there are in case of two triangles every angles equals - so what what result from this - how are the sides corresponding to every angles ?

jhonyy9:

@Kyky232 what you think ?

Kyky232:

I think Im confused

jhonyy9:

above we proven that every angles are equal yes ?

nevaehhh:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @braids36 h2o \(\color{#0cbb34}{\text{End of Quote}}\) w a t e r

Kyky232:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 above we proven that every angles are equal yes ? \(\color{#0cbb34}{\text{End of Quote}}\) yes.

jhonyy9:

so than we know that every corresponding angles are equal from this what result ? not that every corresponding sides are equal too ?

jhonyy9:

ty

Timmyspu:

Or what are they going to add up to in the end?

jhonyy9:

@Kyky232 questions pls ?

Kyky232:

This information would not help to prove the triangles are congruent through SSS. SSS= Side-Side-Side, and we only have one piece of information about sides. The other two are about angles. If we had three sides then it would work, but we don't. This information would help us prove triangles GHI and RST are congruent through ASA or AAS, but not SSS? Eh?